An 10-year corporate bond has a 8.3% coupon rate. What should be the bond's price if the required return is 8% and the bond pays coupons semiannually? (answer in $s- xxxx.xx, with no $ sign needed) Bond Value 1000.00
Answer #1

V = Value of bond = ?

I = Annual interest payable on bond = 83

K = Required rate of return = 8%

RV = Redemption value of bond = 1,000

N = Maturity period of bond = 10 years

V = I/2 x PVAF (k/2 , 2N ) + RV x PVF (k/2 , 2N )

= 83/2 x PVAF ( 4% , 20 ) + 1,000 x PVF ( 4% , 20 )

= 41.5 x 13.59 + 1,000 x 0.456

= 564 + 456

= 1,020

Hence, bonds current price is 1020